What is the probability of randomly selecting a person who has a November birthday? (November has 30 days)
Answer #1

Probability of a person having birthday on any day of the year = 1/365

November has 30 days, so the options multiply by 30

So, probability = 30/365 = 6/73 = 0.0822

[Please give me a Thumbs Up if you are satisfied with my answer. If you are not, please comment on it, so I can edit the answer. Thanks.]